### v1

# Cardinalities in proofreading

Abstract
In 1975, G. Pólya pointed out that if two proofreaders found $a$, resp. $b$, errors in a text, and both found $c$ errors, then the unknown number $e$ of all errors is approximated by $e\approx\frac{ab}{c}$. We devise a realistic combinatorial model of Pólya’s result and estimate its performance by the Probabilistic Method.
1 Introduction. The $(p,n)$-proofreader
In [8], see also [1, 3, 4, 7], G. Pólya devised a conjuring trick. Suppose that a proofreader finds $a$ errors in a text, that another proofreader finds $b$ errors, and that the number of errors detected by both is $c$. How can one guess, with these data, the unknown number $e$ ($>0$) of all errors? Well, $a\approx pe$ and $b\approx qe$, where $p$ and $q$ ($\in(0,1]$) are probabilities of error detection for the proofreaders. Also, $c\approx pqe$ due to their independence. Then
$\frac{ab}{c}\approx\frac{peqe}{pqe}=e\,.$
Pólya’s trick, brilliant as it is, is not a rigorous mathematical theorem. In our article, we elaborate it in a combinatorial theorem. We state this Theorem 1.3 below and prove it in the next section. The last third section contains comments on probabilistic vs. discrete models. For $n$ in $\mathbb{N}$ (=$=\{1,2,\dots\}$) we set $[n]=\{1,2,\dots,n\}$. Let $B_{n}=\{0,1\}^{n}$ be the set of binary $n$-tuples
[…]
We denote by $\|\overline{b}\|=\sum_{i=1}^{n}b_{i}$ the number of $1$s in $\overline{b}$. For a finite set $A$ we denote by $|A|$ ($\in\mathbb{N}_{0}=\{0,1,\dots\}$) its cardinality (number of elements), for example $|B_{n}|=2^{n}$. If $\overline{s}$ is a finite tuple, then $|\overline{s}|$ denotes the number of terms in it. By $\mathbb{Q}$ we denote the set of fractions.
Let $n\in\mathbb{N}$ and $p=\frac{k}{l}\in(0,1]\cap\mathbb{Q}$, with $k,l\in\mathbb{N}$, $k\leq l$ and coprime $k,l$. We always assume that $p>0$. The $(p,n)$-proofreader is the pair $(P,\overline{s})$ in which
$P=P(\overline{a},\,\overline{b})\colon B_{n}\times B_{n}\to B_{n},\ P(\overline{a},\,\overline{b})=(a_{1}b_{1},\,a_{2}b_{2},\,\dots,\,a_{n}b_{n})\,,$
and
$\overline{s}=(\overline{s}_{1},\,\overline{s}_{2},\,\dots,\,\overline{s}_{N}),\ N=l^{n}\,\text{ and }\,\overline{s}_{j}\in B_{n}\,,$
is a uniquely determined $N$-tuple defined as follows. We take the map $f\colon[l]\to\{0,1\}$, $f(x)=1$ for $x\in[k]$ and $f(x)=0$ for $x\in[l]\setminus[k]$, and extend it to
$f\colon[l]^{n}\to B_{n},\ f((a_{1},\,a_{2},\,\dots,\,a_{n}))=(f(a_{1}),\,f(a_{2}),\,\dots,\,f(a_{n}))\,.$
Then we take the lexicographic ordering
$(\overline{u}_{1},\,\overline{u}_{2},\,\dots,\,\overline{u}_{N})$
of $[l]^{n}$, so that $\overline{u}_{1}=(1,1,\dots,1)$, $\overline{u}_{2}=(1,1,\dots,1,2)$, $\dots$, $\overline{u}_{N}=(l,l,\dots,l)$. Any ordering of $[l]^{n}$ would do, but we want that $(p,n)$-proofreaders be unique objects. Finally, we set
$\overline{s}=(\overline{s}_{1},\,\overline{s}_{2},\,\dots,\,\overline{s}_{N})=(f(\overline{u}_{1}),\,f(\overline{u}_{2}),\,\dots,\,f(\overline{u}_{N}))\,.$
Let $(P,\overline{s})$ be the $(p,n)$-proofreader and $\overline{t}\in B_{n}$ be a text, in which the errors are positions of $1$s. Then for $j\in[N]=[l^{n}]$, where $l$ is the denominator of $p$, the proofreader finds in the text $\overline{t}$ in the $j$-th run $\overline{s}_{j}$ exactly $\|P(\overline{s}_{j},\overline{t})\|$ errors. We show that proofreaders work in the desired way.
Proposition 1.1
Let $(P,\overline{s})$, where
$\overline{s}=(\overline{s}_{1},\,\overline{s}_{2},\,\dots,\,\overline{s}_{N})\ \ (\in B_{n}^{N})\,,$
be the $(p,n)$-proofreader. Then for every set $I\subset[n]$ we have
$N^{-1}\cdot\big{|}\big{\{}j\in[N]\colon\;i\in I\Rightarrow s_{j,\,i}=1\big{\}}\big{|}=p^{|I|}\,.$
Proof. Let $\overline{s}$ and $I$ be as stated. It is easy to see from the definition of $\overline{s}$ that the number of $j\in[N]$ such that $s_{j,i}=1$ for every $i\in I$ is $k^{|I|}\cdot l^{n-|I|}$. But then
$N^{-1}\cdot k^{|I|}\cdot l^{n-|I|}=\frac{k^{|I|}l^{n-|I|}}{l^{n}}=\bigg{(}\frac{k}{l}\bigg{)}^{|I|}=p^{|I|}\,.$
$\Box$
Proofreading of a text is usually done only once, or a few times. We think of the $N$ runs $\overline{s}_{j}$ as of performances of the proofreader in possible worlds ([9]).
We need two independent proofreaders. Let $n\in\mathbb{N}$, $p,q\in(0,1]\cap\mathbb{Q}$ and let $(P,\overline{s})$ and $(P,\overline{u})$ be the $(p,n)$-proofreader and the $(q,n)$-proofreader, respectively, with $|\overline{s}|=M$ and $|\overline{u}|=N$. Two (independent) $(p,q,n)$-proofreaders is the pair $(P,\overline{S})$ such that
$\overline{S}$ $=$ $(\overline{S}_{1},\,\overline{S}_{2},\,\dots,\,\overline{S}_{MN})$
$=$ $((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,(\overline{s}_{1},\,\overline{u}_{N}),\,(\overline{s}_{2},\,\overline{u}_{1}),\,(\overline{s}_{2},\,\overline{u}_{2}),\,\dots$
$\dots,\,(\overline{s}_{2},\,\overline{u}_{N}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{1}),\,(\overline{s}_{M},\,\overline{u}_{2})\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))$
($\in(B_{n}^{2})^{MN}$). Again, any ordering of the $MN$ pairs $(\overline{s}_{j},\overline{u}_{j^{\prime}})$ would do, but we want that pairs of proofreaders be unique. For any $j\in[MN]$ we write
$\overline{S}_{j}=\big{(}\overline{s}_{v_{j}},\,\overline{u}_{w_{j}}\big{)}\,.$
We show that two proofreaders work in the desired way.
Proposition 1.2
Let $(P,\overline{S})$, where
$\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,,$
be two $(p,q,n)$-proofreaders. The following holds.
  1. 1.
    For every set $I\subset[n]$ we have
    ${\textstyle\frac{1}{MN}\cdot\big{|}\big{\{}j\in[MN]\colon\;i\in I\Rightarrow s_{v_{j},\,i}=1\big{\}}\big{|}=p^{|I|}\,.}$
  2. 2.
    For every set $I\subset[n]$ we have
    ${\textstyle\frac{1}{MN}\cdot\big{|}\big{\{}j\in[MN]\colon\;i\in I\Rightarrow u_{w_{j},\,i}=1\big{\}}\big{|}=q^{|I|}\,.}$
  3. 3.
    For every set $I\subset[n]$ we have
    ${\textstyle\frac{1}{MN}\cdot\big{|}\big{\{}j\in[MN]\colon\;i\in I\Rightarrow s_{v_{j},\,i}=u_{w_{j},\,i}=1\big{\}}\big{|}=(pq)^{|I|}\,.}$
Proof. Let $\overline{S}$ and $I$ be as stated. 1. Let $r$ be the number of $j\in[M]$ such that $s_{j,i}=1$ for every $i\in I$. By Proposition 1.1, $\frac{r}{M}=p^{|I|}$. Thus the stated product equals
${\textstyle\frac{1}{MN}\cdot rN=\frac{r}{M}=p^{|I|}\,.}$
2. The argument is symmetric, with the number $s$ of $j^{\prime}\in[N]$ such that $u_{j^{\prime},i}=1$ for every $i\in I$. 3. Now the stated product equals
${\textstyle\frac{1}{MN}\cdot rs=\frac{r}{M}\cdot\frac{s}{N}=p^{|I|}\cdot q^{|I|}=(pq)^{|I|}\,.}$
$\Box$
We state our main result, a combinatorial version of Pólya’s trick.
Theorem 1.3
Let $(P,\overline{S})$, where
$\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,,$
be two $(p,q,n)$-proofreaders and let
${\textstyle c(p,\,q)=\frac{1}{2}\big{(}\frac{1}{p}+\frac{1}{q}+\frac{2}{pq}\big{)}^{-1}\,.}$
Then for every text $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$ errors we have for every number $a\in[0,m]$ that
${\textstyle\frac{1}{MN}}\cdot\big{|}\big{\{}j\in[MN]\colon\;\left|\frac{\|P(\overline{s}_{v_{j}},\,\overline{t})\|\cdot\|P(\overline{u}_{w_{j}},\,\overline{t})\|}{\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|}-m\right|\leq a\big{\}}\big{|}\geq$
${\textstyle\geq 1-6\exp(-\frac{2c(p,\,q)^{2}a^{2}}{m})\,.}$
For $\|P(\overline{s}_{v_{j}},P(\overline{u}_{w_{j}},\overline{t}))\|=0$ we define the fraction as +$\infty$+ ∞.
Pólya’s estimate has distance at most $a$ from the number of errors for at least the fraction $1-\dots$ of pairs $(\overline{s}_{v_{j}},\overline{u}_{w_{j}})$ of runs of the two proofreaders. Next, in one extreme regime of Theorem 1.3 we minimize the bound $a$ on the distance. In the other extreme we maximize the fraction of pairs of runs for which the distance is still reasonably small. Deduction of both corollaries is immediate.
Corollary 1.4
In the setup of Theorem 1.3, if $\omega(m)\colon\mathbb{N}\to\mathbb{N}$ is any function such that $\omega(m)\leq\sqrt{m}$ and $\omega(m)\to+\infty$, then for every text $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$ errors we have
${\textstyle\frac{1}{MN}}\cdot\big{|}\big{\{}j\in[MN]\colon\;\left|\frac{\|P(\overline{s}_{v_{j}},\,\overline{t})\|\cdot\|P(\overline{u}_{w_{j}},\,\overline{t})\|}{\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|}-m\right|\leq\omega(m)\sqrt{m}\big{\}}\big{|}\geq$
$\geq 1-6\exp(-2c(p,\,q)^{2}\omega(m)^{2})\to 1\ \ (m\to+\infty)\,.$
Corollary 1.5
In the setup of Theorem 1.3, if $\varepsilon\in(0,1)$ then for every text $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$ errors we have
${\textstyle\frac{1}{MN}}\cdot\big{|}\big{\{}j\in[MN]\colon\;\left|\frac{\|P(\overline{s}_{v_{j}},\,\overline{t})\|\cdot\|P(\overline{u}_{w_{j}},\,\overline{t})\|}{\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|}-m\right|\leq\varepsilon m\big{\}}\big{|}\geq$
$\geq 1-6\exp(-2\varepsilon^{2}c(p,\,q)^{2}m)\to 1\ \ (m\to+\infty)\,.$
We prove Theorem 1.3 next.
2 Proof of Theorem 1.3
We deduce Theorem 1.3 from two propositions.
Proposition 2.1
Suppose that $(P,\overline{S})$, where
$\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,,$
are two $(p,q,n)$-proofreaders and that $a>0$ is a real number. Then for every text $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$ errors we have
${\textstyle\frac{1}{MN}\cdot\big{|}\big{\{}j\in[MN]\colon\;\big{|}\|P(\overline{s}_{v_{j}},\,\overline{t})\|-pm\big{|}\leq a\big{\}}\big{|}\geq 1-2\exp(-\frac{2a^{2}}{m})\,.}$
An analogous bound holds when $\overline{s}_{v_{j}}$ replaced by $\overline{u}_{w_{j}}$, and $pm$ by $qm$.
Proposition 2.2
Suppose that $(P,\overline{S})$, where
$\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,,$
are two $(p,q,n)$-proofreaders and that $a>0$ is a real number. Then for every text $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$ errors we have
${\textstyle\frac{1}{MN}\cdot\big{|}\big{\{}j\in[MN]\colon\;\big{|}\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|-pqm\big{|}\leq a\big{\}}\big{|}\geq 1-2\exp(-\frac{2a^{2}}{m})\,.}$
We prove both propositions at the conclusion of this section. Now we deduce from them Theorem 1.3.
Deduction of Theorem 1.3 from Propositions 2.1 and 2.2. Suppose that $(P,\overline{S})$, where
$\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,,$
are two $(p,q,n)$-proofreaders, that $\overline{t}\in B_{n}$ is a text with $\|\overline{t}\|=m\geq 1$ errors and that $a\in[0,m]$. We set
${\textstyle b=c(p,\,q)\cdot a=\frac{1}{2}\big{(}\frac{1}{p}+\frac{1}{q}+\frac{2}{pq}\big{)}^{-1}\cdot a\ \ (\leq\frac{pqm}{2})\,.}$
For $j\in[MN]$, let $A$ be the set of $j$ such that $|\,\|P(\overline{s}_{v_{j}},\overline{t})\|-pm|\leq b$, $B$ be the set of $j$ such that $|\,\|P(\overline{u}_{w_{j}},\overline{t})\|-qm|\leq b$, $C$ be the set of $j$ such that
$|\,\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|-pqm|\leq b$
and, finally, let $D$ be the set we are interested in, the set of $j$ such that
$|R-m|:=\big{|}\,\|P(\overline{s}_{v_{j}},\,\overline{t})\|\cdot\|P(\overline{u}_{w_{j}},\,\overline{t})\|\cdot\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|^{-1}-m\big{|}\leq a$
($\|\dots\|\cdot\|\dots\|\cdot\|\dots\|^{-1}=+\infty$ if $\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|=0$). Thus if $j\in A$, resp. $j\in B$, then
[…]
where $|\delta_{1}|\leq\frac{b}{pm}$, resp. $|\delta_{2}|\leq\frac{b}{qm}$. Similarly, if $j\in C$ then
${\textstyle\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|=(1+\delta_{3})pqm,\ |\delta_{3}|\leq\frac{b}{pqm}\leq\frac{1}{2}\,.}$
If $j\in A\cap B\cap C$ then
$|R-m|$ $=$ ${\textstyle\big{|}\frac{(1+\delta_{1})(1+\delta_{2})}{1+\delta_{3}}m-m\big{|}}$
$=$ ${\textstyle m\left|(1+\delta_{1}+\delta_{2}+\delta_{1}\delta_{2})\sum_{k=0}^{\infty}(-1)^{k}\delta_{3}^{k}-1\right|}$
$=$ ${\textstyle m\left|\sum_{k=1}^{\infty}(-1)^{k}\delta_{3}^{k}+(\delta_{1}+\delta_{2}+\delta_{1}\delta_{2})\sum_{k=0}^{\infty}(-1)^{k}\delta_{3}^{k}\right|}$
$\leq$ $m\left(2|\delta_{3}|+2(|\delta_{1}|+|\delta_{2}|+|\delta_{1}|\cdot|\delta_{2}|)\right)$
$\leq$ $2b{\textstyle\big{(}\frac{1}{pq}+\frac{1}{p}+\frac{1}{q}+\frac{1}{pq}\big{)}=a\,,}$
and hence $D\supset A\cap B\cap C$. By Propositions 2.1 and 2.2,
${\textstyle\frac{1}{MN}\cdot|D|\geq 1-\frac{|A^{c}|+|B^{c}|+|C^{c}|}{MN}\geq 1-6\exp(-\frac{2b^{2}}{m})=1-6\exp(-\frac{2c(p,\,q)^{2}a^{2}}{m})\,.}$
$\Box$
We develop tools for proofs of Propositions 2.1 and 2.2. Let $(P,\overline{S})$, where
$\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,,$
be two $(p,q,n)$-proofreaders. We consider the finite uniform probability space
$P=P(MN)=([MN],\,\mathcal{P}([MN]),\,\mathrm{Pr})\,,$
in which $\mathrm{Pr}(\{j\})=\frac{1}{MN}$, and for $i\in[n]$ the three $n$-tuples of $P$-events
$A_{i}=\{j\in[MN]\colon\;s_{v_{j},i}=1\},\ A_{i}^{\prime}=\{j\in[MN]\colon\;u_{w_{j},i}=1\}$
and
$C_{i}=\{j\in[MN]\colon\;s_{v_{j},i}=u_{w_{j},i}=1\}\,.$
By Propositions 1.1 and 1.2, each of these three $n$-tuples is independent. Further, $\mathrm{Pr}(A_{i})=p$, $\mathrm{Pr}(A_{i}^{\prime})=q$ and $\mathrm{Pr}(C_{i})=pq$ for every $i\in[n]$. Recall that $B^{c}$ denotes the complementary event to the event $B$ in a probability space. For example,
$A_{i}^{c}=[MN]\setminus A_{i}=\{j\in[MN]\colon\;s_{v_{j},i}=0\}\,.$
Let $X\subset[n]$ be a nonempty set. For any set $Y\subset X$ we consider the three $P$-events
${\textstyle A_{Y,X}=\bigcap_{i\in Y}A_{i}\cap\bigcap_{i\in X\setminus Y}A_{i}^{c},\ A_{Y,X}^{\prime}=\bigcap_{i\in Y}A_{i}^{\prime}\cap\bigcap_{i\in X\setminus Y}(A_{i}^{\prime})^{c}}$
and
${\textstyle C_{Y,X}=\bigcap_{i\in Y}C_{i}\cap\bigcap_{i\in X\setminus Y}C_{i}^{c}\,.}$
We use the following key theorem which restates [2, Corollary A.1.7 on p. 325]. It is proven on [2, pp. 321–325].
Theorem 2.3
Let $n\in\mathbb{N}$, $X\subset[n]$ with $|X|=m\geq 1$ and let $a>0$ be a real number. With the above notation we have the bound
${\textstyle\sum_{\begin{subarray}{c}Y,\,Y\subset X\\ |\,|Y|-pm\,|\leq a\end{subarray}}\mathrm{Pr}(A_{Y,X})\geq 1-2\exp(-\frac{2a^{2}}{m})\,.}$
Two analogous bounds hold when $A_{Y,X}$ is replaced by $A_{Y,X}^{\prime}$, resp. $C_{Y,X}$, and $pm$ by $qm$, resp. $pqm$.
We explain in detail this restatement. Let $m\in\mathbb{N}$, $p\in[0,1]$, $X_{i}$ for $i\in[m]$ be $m$ independent two-valued real random variables in a probability space, with the common distribution $\mathrm{Pr}(X_{i}=1-p)=p$ and $\mathrm{Pr}(X_{i}=-p)=1-p$, and let
$X_{0}=X_{1}+X_{2}+\dots+X_{m}\,.$
The random variable $X_{0}$ has the centered binomial distribution $B(m,p)-pm$. The estimate [2, Corollary A.1.7], more precisely a particular case of it with all $p_{i}=p$, says the following.
Theorem 2.4 (Corollary A.1.7 in [2])
For every real number $a>0$,
${\textstyle\mathrm{Pr}(|X_{0}|>a)<\exp(-\frac{2a^{2}}{m})\,.}$
What does the independence of random variables $X_{i}$, $i\in[m]$, which is assumed in Theorem 2.4, precisely mean? Some events $C_{1}$, $C_{2}$, $\dots$, $C_{m}$ in a probability space are independent, if for every set $I\subset[m]$ we have
${\textstyle\mathrm{Pr}\big{(}\bigcap_{i\in I}C_{i}\big{)}=\prod_{i\in I}\mathrm{Pr}(C_{i})\,.}$
These events are strongly independent, if every $m$-tuple of events $D_{1}$, $D_{2}$, $\dots$, $D_{m}$, where $D_{i}\in\{C_{i},C_{i}^{c}\}$, is independent. A well known lemma, whose proof we leave for the interested reader as an exercise, says that any independent events are strongly independent. Discrete (i.e., with at most countable images) real random variables $Y_{1}$, $Y_{2}$, $\dots$, $Y_{m}$ are independent if for every $m$-tuple
$\overline{a}=(a_{1},\,a_{2},\,\dots,\,a_{m})\in\mathbb{R}^{m}$
the events $\mathrm{Pr}(Y_{i}=a_{i})$, $i\in[m]$, are independent. If the $Y_{i}$ are two-valued, the image of $Y_{i}$ being $\{a_{i},b_{i}\}$ with $a_{i}\neq b_{i}$, then, due to the mentioned lemma, the random variables $Y_{i}$ are independent if for any single $m$-tuple $\overline{c}\in\mathbb{R}^{m}$ with $c_{i}\in\{a_{i},b_{i}\}$ the events $\mathrm{Pr}(Y_{i}=c_{i})$, $i\in[m]$, are independent.
We deduce Theorem 2.3 from Theorem 2.4. We consider only events $A_{Y,X}$, for events $A_{Y,X}^{\prime}$ and $C_{Y,X}$ we would argue similarly, relying on part 2 and 3 of Proposition 2.2. So let $(P,\overline{S})$, with
$\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,,$
be as above two $(p,q,n)$-proofreaders and let $X\subset[n]$ with $|X|=m\geq 1$. We may assume that $X=[m]$. For $i\in[m]$ we define two-valued $P(MN)$-random variables $X_{i}$ by setting $X_{i}(j)=1-p$ if $j\in A_{i}$, and $X_{i}(j)=-p$ if $j\in[MN]\setminus A_{i}$. We set $X_{0}=\sum_{i=1}^{m}X_{i}$. Then for any subset $Y\subset X$ we have
$\mathrm{Pr}(A_{Y,X})=\mathrm{Pr}(X_{0}=|Y|-pm)\,.$
By part 1 of Proposition 2.2 and the above discussion of independence, the random variables $X_{i}$, $i\in[m]$, are independent. Let $a>0$ be a real number. By applying Theorem 2.4 to the random variable $X_{0}$ we get the estimate in Theorem 2.3.
We proceed to proofs of Propositions 2.1 and 2.2.
Proof of Proposition 2.1. Suppose that $(P,\overline{S})$, where
$\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,,$
are two $(p,q,n)$-proofreaders and that $a>0$ is a real number. Let $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$. We may assume that the $1$s of $\overline{t}$ are on the first $m$ positions. We consider the $P(MN)$-events $A_{i}$, $i\in[m]$. Let $Y\subset[m]=X$. Then
${\textstyle\frac{1}{MN}|\{j\in[MN]\colon\;\text{for $i\in[m]$ we have $s_{v_{j},i}=1$ iff $i\in Y$}\}|=\mathrm{Pr}(A_{Y,X})\,.}$
Thus for any $k\in\mathbb{N}_{0}$ we have
${\textstyle\frac{1}{MN}\cdot\big{|}\big{\{}j\in[MN]\colon\;\|P(\overline{s}_{v_{j}},\,\overline{t})\|=k\big{\}}\big{|}=\sum_{\begin{subarray}{c}Y,\,Y\subset X\\ |Y|=k\end{subarray}}\mathrm{Pr}(A_{Y,X})\,.}$
The stated estimate therefore follows from the first part of Theorem 2.3. The second part of Proposition 2.1 follows by a very similar argument from the second part of Theorem 2.3. $\Box$
Proof of Proposition 2.2. Let $(P,\overline{S})$, $a>0$ and $\overline{t}$ be as in the previous proof. We consider the $P(MN)$-events $C_{i}$, $i\in[m]$, and argue as in the previous proof, relying on the third part of Theorem 2.3. $\Box$
3 Concluding remarks
By [5, p. 109], “The subject of probability is wholly free from real paradoxes.” This is not precise — not only books like [10] are devoted to paradoxes in probability theory, but some of these paradoxes are truly real. At their roots is the unrealistic nature, or even total absence, of physical experiments that should be modeled by the probabilistic model, or PM, at hand. In these paradoxical cases we often have a PM that does not model anything realistic. For example, take well known Bertrand’s paradox which asks about the probability that random chord in the unit circle in the plane has length at least $\sqrt{3}$ (the side of the inscribed equilateral triangle). The noun phrase “random chord” is problematic because it does not refer to anything clearly defined. In contrast, noun phrases like “random (finite!) graph” or “random subset (of a finite set)” are unproblematic; these concepts are easily modeled by finite uniform probability spaces and events in them, like the above $P(MN)$ and the $P(MN)$-events $A_{i}$, $A_{i}^{\prime}$ and $C_{i}$, $i\in[n]$. Mathematical chords have zero width and form an uncountable set. People try to resolve Bertrand’s paradox by rushing in defining PM on the set of mathematical chords. Instead, they should step back and consider physical chords with width $w>0$. A realistic and convincing physical discrete model, or DM, for selection of random physical chords should be devised. Only when such DM is understood, we can send $w$ to 00 and see what happens. This approach is implemented in [6]. We hope to rework [6] in near future in a better form.
In this article we approached Pólya’s proofreaders from the DM perspective. Returning to Proposition 1.1, we see that if $(P,\overline{s})$, where
$\overline{s}=(\overline{s}_{1},\,\overline{s}_{2},\,\dots,\,\overline{s}_{N})\,,$
is the $(p,n)$-proofreader, then in the uniform probability space
$P(N)=([N],\,\mathcal{P}([N]),\,\mathrm{Pr})\,,$
with $\mathrm{Pr}(\{j\})=\frac{1}{N}$, the $n$ $P(N)$-events
$A_{i}=\{j\in[N]\colon\;s_{j,i}=1\},\ i\in[n]\,,$
are independent and model random subset of $[n]$ with probability of an element equal to $p$. However, our construction requires $p$ ($\in(0,1]$) be a rational number. This is just as right because we cannot have a realistic physical experiment with probabilities like $\frac{1}{\sqrt{2}}$. Thus we conclude our article with the comment, which may come as a surprise to some, that the random graph $G(n,p)$ with probability of an edge equal to an irrational number $p$ is not really a realistic model.