### v2

# Cardinalities in proofreading

Abstract
In 1975, G. Pólya suggested that if two proofreaders found $a$ and $b$ errors in a text, of which $c$ errors were found by both of them, then a reasonable approximation of the unknown number $e$ of all errors is $e\approx ab/c$. We justify this formula by constructing a realistic model of proofreaders and estimating the efficiency of this model with the help of the bound on large deviations in the Probabilistic Method. In conclusion we discuss the distinction between realistic and probabilistic models of problems.
Dedicated to the memory of my mother Blanka Klazarová (1937–2026)
1 Introduction. $(p,n)$-proofreaders
In [7], G. Pólya considered the following problem (see also [1, 3, 4, 6]). Suppose that a proofreader found $a$ errors in a text, that another proofreader found $b$ errors, and that the number of errors detected by both of them is $c$. With these data, how could one approximate the unknown number $e$ of all errors in the text? Pólya’s solution is this. We have $a\approx pe$ and $b\approx qe$, where $p$ and $q$ in $(0,1]$ are the probabilities of error detection for the two proofreaders. Their independence gives $c\approx pqe$. Thus we get the approximation
$e=\frac{peqe}{pqe}\approx\frac{ab}{c}\,.$
Pólya’s interesting argument is not supported by any rigorous theorem. In this article we provide one. It is Theorem 1.5 which we prove in Section 2. In Section 3 we discuss the two kinds of mathematical models one can devise to solve problems: realistic models and probabilistic models.
For $n$ in $\mathbb{N}=\{1,2,\dots\}$ we set $n]=\{1,2,\dots,n\}$. Let $B_{n}=\{0,1\}^{n}$ be the set of binary $n$-tuples
$\overline{b}=(\overline{b}_{1},\,\overline{b}_{2},\,\dots,\,\overline{b}_{n}),\ \overline{b}_{i}\in\{0,\,1\}\,.$
We denote by
$\|\overline{b}\|=\sum_{i=1}^{n}\overline{b}_{i}$}
the number of $1$s in $\overline{b}$. For a finite set $A$ we denote by $|A|$ ($\in\mathbb{N}_{0}=\{0,1,\dots\}$) the number of its elements. For example, $|B_{n}|=2^{n}$. For a tuple $\overline{s}$ we denote by $|\overline{s}|$ its length, the number of terms in $\overline{s}$. By $\mathbb{Q}$ we denote the set of fractions.
Let $p=k/l$ be in $(0,1]\cap\mathbb{Q}$, with $\in\mathbb{N}, klk\leq l$ and coprime $k,l$, and let $n\in\mathbb{N}$. We define the key discrete structure of our article.
Definition 1.1 ($(p,n)$-proofreaders)
$(p,n)$-proofreader $\mathcal{P}$ is the pair $\mathcal{P}=(P,\overline{s})$ with the following components which are uniquely determined by $p$ and $n$.
  1. 1.
    $P=P(\overline{a},\overline{b})\colon B_{n}\times B_{n}\to B_{n}$ is a map given by
    $P(\overline{a},\,\overline{b})=(\overline{a}_{1}\overline{b}_{1},\,\overline{a}_{2}\overline{b}_{2},\,\dots,\,\overline{a}_{n}\overline{b}_{n})\,.$
  2. 2.
    The $N$-tuple of $n$-tuples
    $\overline{s}=(\overline{s}_{1},\,\overline{s}_{2},\,\dots,\,\overline{s}_{N})\,$,
    where $\overline{s}_{j}\in B_{n}$ and $N=l^{n}$, is defined as follows. We begin with the map $f\colon[l]\to\{0,1\}$ given by $f(x)=1$ for $x\in[k]$ and $f(x)=0$ for $x\in[l]\setminus[k]$. We extend it to the map $f\colon[l]^{n}\to B_{n}$ by
    $f(a_{1},\,a_{2},\,\dots,\,a_{n})=(f(a_{1}),\,f(a_{2}),\,\dots,\,f(a_{n}))\,.$
    We take the lexicographic ordering $\overline{u}_{1}$, $\overline{u}_{2}$, $\dots$, $\overline{u}_{N}$ of $l]^{n}$, so that $\overline{u}_{1}=(1,1,\dots,1)$, $\overline{u}_{2}=(1,1,\dots,1,2)$, $\dots$, $\overline{u}_{N}=(l,l,\dots,l)$. Finally, we set
    $\overline{s}=(\overline{s}_{1},\,\overline{s}_{2},\,\dots,\,\overline{s}_{N})=(f(\overline{u}_{1}),\,f(\overline{u}_{2}),\,\dots,\,f(\overline{u}_{N}))\,.$
    We call the $N$-tuple $\overline{s}$ the runs of $\mathcal{P}$.
Any ordering of $l]^{n}$ would do. We chose the lexicographic one just concreteness sake.
We show that the proofreaders introduced in Definition 1.1 work according to the parameters $p$ and $n$. The parameter $p$ is the efficiency (probability of error detection) of the proofreader, and the parameter $n$ is the length of the text submitted for proofreading. $\mathcal{P}=(P,\overline{s})$ works in $N$ runs $\overline{s}_{j}$, $j\in[N]$. If $(\overline{s}_{j})_{i}=1$ then in the $j$-th run $\mathcal{P}$ pays attention to the $i$-th position in the text, and if $(\overline{s}_{j})_{i}=0$ then $\mathcal{P}$ ignores it. So if a text $\overline{t}\in B_{n}$ (in which $1$s are errors) is submitted for proofreading, then $\mathcal{P}$ in the run $\overline{s}_{j}$ detects
$\|P(\overline{s}_{j},\,\overline{t})\|\ \ (\in\mathbb{N}_{0})$
errors in $\overline{t}$. The next proposition shows that for every set $I\subset[n]$ of positions in the text, $\mathcal{P}$ pays attention to all positions in II in exactly $p^{|I|}\cdot N$ of the $N$ runs.
Proposition 1.2
Let $\overline{s}=(\overline{s}_{1},\overline{s}_{2},\dots,\overline{s}_{N})$ be the runs of the $(p,n)$-proofreader. Then for every set $I\subset[n]$ we have
$N^{-1}\cdot\big|\big\{j\in[N]\colon\;\text{$(\overline{s}_{j})_{i}=1$ for every $i\in I$}\}\big|=p^{|I|}\,.$
Proof. Let $\overline{s}$, $p$, $n$, and $I$ be as stated. It is easy to see from the definition of $\overline{s}$ that the number of indices $j\in[N]$ such that $(\overline{s}_{j})_{i}=1$ for every $i\in I$ equals
$k^{|I|}\cdot l^{n-|I|}$
because it is the number of words of length $n$ over the alphabet $[l]$ that have at every position $i\in I$ a letter from $[k]$ ($\subset[l]$) . But then
$N^{-1}\cdot k^{|I|}\cdot l^{n-|I|}=\frac{k^{|I|}\cdot l^{n-|I|}}{l^{n}}=\bigg(\frac{k}{l}\bigg)^{|I|}=p^{|I|}\,.$
$\Box$
For Pólya’s argument we need two independent proofreaders. Suppose that $p,q\in(0,1]\cap\mathbb{Q}$ and that $n\in\mathbb{N}$.
Definition 1.3 (two $(p,q,n)$-proofreaders)
Two $(p,q,n)$-proofreaders is the pair $\mathcal{R}=(P,\overline{S})$ with the following components which are uniquely determined by $p$, $q$ and $n$.
  1. 1.
    The map $P$ is the same as in Definition 1.1.
  2. 2.
    We construct the $MN$-tuple $\overline{S}$ of pairs of $n$-tuples as follows. Let $(P,\overline{s})$ be the $(p,n)$-proofreader, $(P,\overline{u})$ be the $(q,n)$-proofreader, and let $M=|\overline{s}|$ and $N=|\overline{u}|$. We set
    $\overline{S}$ $=$ $(\overline{S}_{1},\,\overline{S}_{2},\,\dots,\,\overline{S}_{MN})$
    $=$ $((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,(\overline{s}_{1},\,\overline{u}_{N}),\,(\overline{s}_{2},\,\overline{u}_{1}),\,(\overline{s}_{2},\,\overline{u}_{2}),\,\dots,\,(\overline{s}_{2},\,\overline{u}_{N})\,$,
    $\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{1}),\,(\overline{s}_{M},\,\overline{u}_{2}),\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\ \ (\in(B_{n}\times B_{n})^{MN})\,.$
    We call the $MN$-tuple $\overline{S}$ the runs of $\mathcal{R}$.
Any ordering of the $MN$ pairs $(\overline{s}_{j},\overline{u}_{j^{\prime}})$ would do, but we again want to have a uniquely defined object. For any $j\in[MN]$ we write
$\overline{S}_{j}=\big(\overline{s}_{v_{j}},\,\overline{u}_{w_{j}}\big)\,.$
In the next proposition, we show that for every set $I\subset[n]$ of positions in the text submitted for proofreading, the two proofreaders pay attention to these positions mutually independently.
Proposition 1.4
Let $\overline{S}=((\overline{s}_{1},\overline{u}_{1}),(\overline{s}_{1},\overline{u}_{2}),\dots,\dots,(\overline{s}_{M},\,\overline{u}_{N}))$ be the runs of two $(p,q,n)$-proofreaders. The following holds.
  1. 1.
    For every set $I\subset[n]$ we have
    $\frac{1}{MN}\cdot\big|\big\{j\in[MN]\colon\;\text{$(\overline{s}_{v_{j}})_{i}=1$ for every $i\in I$}\big\}\big|=p^{|I|}\,.$}
  2. 2.
    For every set $I\subset[n]$ we have
    $\frac{1}{MN}\cdot\big|\big\{j\in[MN]\colon\;\text{$(\overline{u}_{w_{j}})_{i}=1$ for every $i\in I$}\big\}\big|=q^{|I|}\,.$}
  3. 3.
    For every set $I\subset[n]$ we have
    $\frac{1}{MN}\cdot\big|\big\{j\in[MN]\colon\;\text{$(\overline{s}_{v_{j}})_{i}=(\overline{u}_{w_{j}})_{i}=1$ for every $i\in I$}\big\}\big|=(pq)^{|I|}\,.$}
Proof. Let $\overline{S}$ and $I$ be as stated.
1. Let $r$ be the number of indices $j\in[M]$ such that $(\overline{s}_{j})_{i}=1$ for every $i\in I$. By Proposition 1.2, $\frac{r}{M}=p^{|I|}$. Thus the displayed product equals
$\frac{1}{MN}\cdot rN=\frac{r}{M}=p^{|I|}\,.$}
2. The argument is symmetric, with the number $s$ of indices $j\in[N]$ such that $(\overline{u}_{j})_{i}=1$ for every $i\in I$.
3. Now the displayed product equals
$\frac{1}{MN}\cdot rs=\frac{r}{M}\cdot\frac{s}{N}=p^{|I|}\cdot q^{|I|}=(pq)^{|I|}\,.$}
$\Box$
The following theorem is one of our main results.
Theorem 1.5
Let $\mathcal{R}=(P,\overline{S})$, where
$\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,$,
be two $(p,q,n)$-proofreaders and let
$c(p,\,q)=\frac{1}{2}\big(\frac{1}{p}+\frac{1}{q}+\frac{2}{pq}\big)^{-1}\ \ (\in(0,\,\frac{1}{8}])\,.$}
Then for every text $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$ errors and for every real number $a\in[0,m]$ we have
${\textstyle\frac{1}{MN}}\cdot\big|\big\{j\in[MN]\colon\;\left|\frac{\|P(\overline{s}_{v_{j}},\,\overline{t})\|\cdot\|P(\overline{u}_{w_{j}},\,\overline{t})\|}{\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|}-m\right|\leq a\big\}\big|\geq$
${\textstyle\geq 1-6\exp\big(-\frac{2c(p,\,q)^{2}a^{2}}{m}\big)\,,}$
where for $\|P(\overline{s}_{v_{j}},P(\overline{u}_{w_{j}},\overline{t}))\|=0$ we define the fraction $\frac{\cdots}{0}$ as $\infty$.
The theorem says that the distance between Pólya’s approximation and the actual number $m$ of errors is at most $a$ for at least the displayed fraction 11-$\dots$ of the $MN$ runs of $\mathcal{R}$.
The theorem has two corollaries. We minimize the distance of the approximation from the actual number of errors. Then we maximize the fraction of runs for which the distance is still reasonably small. The proofs are immediate, and we omit them.
Corollary 1.6
In the setup of Theorem 1.5, if $\omega\colon\mathbb{N}\to\mathbb{N}$ is a function such that
[…]
then for every text $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$ errors we have
${\textstyle\frac{1}{MN}}\cdot\big|\big\{j\in[MN]\colon\;\left|\frac{\|P(\overline{s}_{v_{j}},\,\overline{t})\|\cdot\|P(\overline{u}_{w_{j}},\,\overline{t})\|}{\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|}-m\right|\leq\omega(m)\sqrt{m}\big\}\big|\geq$
$\geq 1-6\exp(-2c(p,\,q)^{2}\omega(m)^{2})\to 1\ \ (m\to\infty)\,.$
Corollary 1.7
In the setup of Theorem 1.5, if $\varepsilon\in(0,1)$ then for every text $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$ errors we have
${\textstyle\frac{1}{MN}}\cdot\big|\big\{j\in[MN]\colon\;\left|\frac{\|P(\overline{s}_{v_{j}},\,\overline{t})\|\cdot\|P(\overline{u}_{w_{j}},\,\overline{t})\|}{\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|}-m\right|\leq\varepsilon m\big\}\big|\geq$
$\geq 1-6\exp(-2\varepsilon^{2}c(p,\,q)^{2}m)\to 1\ \ (m\to\infty)\,.$
We prove Theorem 1.5 in the next section.
2 Proof of Theorem 1.5
We deduce Theorem 1.5 from the next proposition.
Proposition 2.1
Suppose that $(P,\overline{S})$, where
$\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,$,
are the two $(p,q,n)$-proofreaders and that $a>0$ is a real number. Then for every text $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$ errors the following holds.
  1. 1.
    We have
    $\frac{1}{MN}\cdot\big|\big\{j\in[MN]\colon\;\big|\|P(\overline{s}_{v_{j}},\,\overline{t})\|-pm\big|\leq a\big\}\big|\geq 1-2\exp(-\frac{2a^{2}}{m})\,.$}
  2. 2.
    We have
    $\frac{1}{MN}\cdot\big|\big\{j\in[MN]\colon\;\big|\|P(\overline{u}_{w_{j}},\,\overline{t})\|-qm\big|\leq a\big\}\big|\geq 1-2\exp(-\frac{2a^{2}}{m})\,.$}
  3. 3.
    We have
    ${\textstyle\frac{1}{MN}\cdot\big|\big\{j\in[MN]\colon\;\big|\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|-pqm\big|\leq a\big\}\big|}$
    ${\textstyle\geq 1-2\exp(-\frac{2a^{2}}{m})\,.}$
We prove it at the end of this section. Now we deduce Theorem 1.5.
Deduction of Theorem 1.5 from Proposition 2.1. Suppose that $(P,\overline{S})$, where
$\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,$,
are two $(p,q,n)$-proofreaders, that $\overline{t}\in B_{n}$ is a text with $\|\overline{t}\|=m\geq 1$ errors and that $a\in[0,m]$ is a real number. We set
$b=c(p,\,q)\cdot a=\frac{1}{2}\big(\frac{1}{p}+\frac{1}{q}+\frac{2}{pq}\big)^{-1}\cdot a\ \ (\leq\frac{pqm}{2})\,.$}
Let $A$ be the set of $j$ ($\in[MN])$ such that $|\,\|P(\overline{s}_{v_{j}},\overline{t})\|-pm|\leq b$, $B$ be the set of $j$ such that $|\,\|P(\overline{u}_{w_{j}},\overline{t})\|-qm|\leq b$, $C$ be the set of $j$ such that
$|\,\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|-pqm|\leq b$
and, finally, let $D$ be the set we are interested in, the set of $j$ such that
$|R-m|:=\big|\,\|P(\overline{s}_{v_{j}},\,\overline{t})\|\cdot\|P(\overline{u}_{w_{j}},\,\overline{t})\|\cdot\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|^{-1}-m\big|\leq a\,$,
where $\|\dots\|\cdot\|\dots\|\cdot\|\dots\|^{-1}=+\infty$ if $\|P(\overline{s}_{v_{j}},P(\overline{u}_{w_{j}},\overline{t}))\|=0$. Thus if $j\in A$, resp. $j\in B$, then
$\text{$\|P(\overline{s}_{v_{j}},\,\overline{t})\|=(1+\delta_{1})pm$, resp. $\|P(\overline{u}_{w_{j}},\,\overline{t})\|=(1+\delta_{2})qm$}\,$,
where $|\delta_{1}|\leq\frac{b}{pm}$, resp. $|\delta_{2}|\leq\frac{b}{qm}$. Similarly, if $j\in C$ then
$\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|=(1+\delta_{3})pqm,\ |\delta_{3}|\leq\frac{b}{pqm}\leq\frac{1}{2}\,.$}
If $j\in A\cap B\cap C$ then
$|R-m|$ $=$ ${\textstyle\big|\frac{(1+\delta_{1})(1+\delta_{2})}{1+\delta_{3}}m-m\big|}$
$=$ ${\textstyle m\left|(1+\delta_{1}+\delta_{2}+\delta_{1}\delta_{2})\sum_{k=0}^{\infty}(-1)^{k}\delta_{3}^{k}-1\right|}$
$=$ ${\textstyle m\left|\sum_{k=1}^{\infty}(-1)^{k}\delta_{3}^{k}+(\delta_{1}+\delta_{2}+\delta_{1}\delta_{2})\sum_{k=0}^{\infty}(-1)^{k}\delta_{3}^{k}\right|}$
$\leq$ $m\left(2|\delta_{3}|+2(|\delta_{1}|+|\delta_{2}|+|\delta_{1}|\cdot|\delta_{2}|)\right)$
$\leq$ $2b{\textstyle\big(\frac{1}{pq}+\frac{1}{p}+\frac{1}{q}+\frac{1}{pq}\big)=a}$
and therefore $D\supset A\cap B\cap C$. By items 1–3 of Proposition 2.1,
$\frac{1}{MN}\cdot|D|\geq 1-\frac{|A^{c}|+|B^{c}|+|C^{c}|}{MN}\geq 1-6\exp(-\frac{2b^{2}}{m})=1-6\exp(-\frac{2c(p,\,q)^{2}a^{2}}{m})$}
(for $X\subset[MN]$ we write $X^{c}$ for the complement $MN]\setminus X$). $\Box$
We develop tools for the proof of Proposition 2.1. Let $(P,\overline{S})$, where
$\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,$,
be two $(p,q,n)$-proofreaders. We consider the finite uniform probability space
$\mathrm{PS}(MN)=([MN],\,\mathcal{P}([MN]),\,\mathrm{Pr})\,$,
where $\mathrm{Pr}(\{j\})=\frac{1}{MN}$, and for $i\in[n]$ the three $n$-tuples of $\mathrm{PS}(MN)$-events
$A_{i}=\{j\in[MN]\colon\;(\overline{s}_{v_{j}})_{i}=1\},\ A_{i}^{\prime}=\{j\in[MN]\colon\;(\overline{u}_{w_{j}})_{i}=1\}$
and
$C_{i}=\{j\in[MN]\colon\;(\overline{s}_{v_{j}})_{i}=(\overline{u}_{w_{j}})_{i}=1\}\,.$
By items 1–3 of Proposition 1.4, each of these three $n$-tuples consists of independent events. Also, $\mathrm{Pr}(A_{i})=p$, $\mathrm{Pr}(A_{i}^{\prime})=q$ and $\mathrm{Pr}(C_{i})=pq$ for every $i\in[n]$. Recall that $B^{c}$ denotes the complementary event to the event $B$. For example,
$A_{i}^{c}=[MN]\setminus A_{i}=\{j\in[MN]\colon\;(\overline{s}_{v_{j}})_{i}=0\}\,.$
Let $X\subset[n]$ be a nonempty set. For any set $Y\subset X$ we consider the three $\mathrm{PS}(MN)$-events
$A_{Y,X}=\bigcap_{i\in Y}A_{i}\cap\bigcap_{i\in X\setminus Y}A_{i}^{c},\ A_{Y,X}^{\prime}=\bigcap_{i\in Y}A_{i}^{\prime}\cap\bigcap_{i\in X\setminus Y}(A_{i}^{\prime})^{c}$}
and
$C_{Y,X}=\bigcap_{i\in Y}C_{i}\cap\bigcap_{i\in X\setminus Y}C_{i}^{c}\,.$}
We employ the following theorem which follows from [2, Corollary A.1.7 on p. 325] which is proven in [2, pp. 321–325].
Theorem 2.2
Let $n\in\mathbb{N}$, $X\subset[n]$ with $|X|=m\geq 1$ and let $a>0$ be a real number. With the above notation we have the following bounds.
  1. 1.
    $\sum_{\begin{subarray}{c}Y,\,Y\subset X\\ |\,|Y|-pm\,|\leq a\end{subarray}}\mathrm{Pr}(A_{Y,X})\geq 1-2\exp(-\frac{2a^{2}}{m})\,.$}
  2. 2.
    $\sum_{\begin{subarray}{c}Y,\,Y\subset X\\ |\,|Y|-qm\,|\leq a\end{subarray}}\mathrm{Pr}(A_{Y,X}^{\prime})\geq 1-2\exp(-\frac{2a^{2}}{m})\,.$}
  3. 3.
    $\sum_{\begin{subarray}{c}Y,\,Y\subset X\\ |\,|Y|-pqm\,|\leq a\end{subarray}}\mathrm{Pr}(C_{Y,X})\geq 1-2\exp(-\frac{2a^{2}}{m})\,.$}
We explain in detail the deduction of this theorem from [2, Corollary A.1.7]. Let $m\in\mathbb{N}$, $p\in[0,1]$, $X_{i}$ for $i\in[m]$ be $m$ independent two-valued real random variables in a probability space, with the common distribution $\mathrm{Pr}(X_{i}=1-p)=p$ and $\mathrm{Pr}(X_{i}=-p)=1-p$, and let
$X_{0}=X_{1}+X_{2}+\dots+X_{m}\,.$
The random variable $X_{0}$ has the centered binomial distribution B(m,p)pmB(m,p)-pm. The estimate [2, Corollary A.1.7], more precisely a particular case of it with all $p_{i}=p$, says the following.
Theorem 2.3 (Corollary A.1.7 in [2])
With the above notation we have for every real number $a>0$ that
$\mathrm{Pr}(|X_{0}|>a)<\exp(-\frac{2a^{2}}{m})\,.$}
What does the independence of random variables $X_{i}$, $i\in[m]$, which is needed in Theorem 2.3, precisely mean? Some events $C_{1}$, $C_{2}$, $\dots$, $C_{m}$ in a probability space are independent, if for every set $I\subset[m]$ we have
$\mathrm{Pr}\big(\bigcap_{i\in I}C_{i}\big)=\prod_{i\in I}\mathrm{Pr}(C_{i})\,.$}
These events are strongly independent, if every $m$-tuple of events $D_{1}$, $D_{2}$, $\dots$, $D_{m}$, where each $D_{i}\in\{C_{i},C_{i}^{c}\}$, is independent. A well known lemma, whose proof we leave for the interested reader as an exercise, says that any independent events are strongly independent. Discrete (i.e., with at most countable images) real random variables $Y_{1}$, $Y_{2}$, $\dots$, $Y_{m}$ are independent if for every $m$-tuple
$\overline{a}=(\overline{a}_{1},\,\overline{a}_{2},\,\dots,\,\overline{a}_{m})\in\mathbb{R}^{m}\,$,
the events $\mathrm{Pr}(Y_{i}=\overline{a}_{i})$, $i\in[m]$, are independent. If the $Y_{i}$ are two-valued, the image of $Y_{i}$ being $\{a_{i},b_{i}\}$ with $a_{i}\neq b_{i}$, then, due to the mentioned lemma, the random variables $Y_{i}$ are independent if for any single $m$-tuple $\overline{c}\in\mathbb{R}^{m}$ with $\overline{c}_{i}\in\{a_{i},b_{i}\}$ the events $\mathrm{Pr}(Y_{i}=c_{i})$, $i\in[m]$, are independent.
We deduce Theorem 2.2 from Theorem 2.3. We consider only item 1, the arguments for items 2 and 3 are similar. So let $(P,\overline{S})$, with
$\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,$,
be as above two $(p,q,n)$-proofreaders and let $X\subset[n]$ with $|X|=m\geq 1$. We may assume that $X=[m]$. For $i\in[m]$ we define two-valued $\mathrm{PS}(MN)$-random variables $X_{i}$ by setting ($j\in[MN]$)
$\text{$X_{i}(j)=1-p$ if $j\in A_{i}$, and $X_{i}(j)=-p$ if $j\in[MN]\setminus A_{i}$}\,.$
We set $X_{0}=\sum_{i=1}^{m}X_{i}$. Then for any subset $Y\subset X$ we have
$\mathrm{Pr}(A_{Y,X})=\mathrm{Pr}(X_{0}=|Y|-pm)\,.$
By item 1 of Proposition 1.4 and the above discussion of independence, the random variables $X_{i}$, $i\in[m]$, are independent. Let $a>0$ be a real number. Applying Theorem 2.3 to the random variable $X_{0}$ we get the estimate in item 1 of Theorem 2.2.
We proceed to the proof of Proposition 2.1.
Proof of Proposition 2.1. We consider only item 1, the arguments for items 2 and 3 are similar. Suppose that $(P,\overline{S})$, where
$\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,$,
are two $(p,q,n)$-proofreaders and that $a>0$ is a real number. Let $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$. We may assume that the $1$s of $\overline{t}$ are on the first $m$ positions. We consider the $\mathrm{PS}(MN)$-events $A_{i}$, $i\in[m]$. Let $Y\subset[m]=X$. Then
$\frac{1}{MN}|\{j\in[MN]\colon\;\text{for $i\in[m]$ we have $(\overline{s}_{v_{j}})_{i}=1$ iff $i\in Y$}\}|=\mathrm{Pr}(A_{Y,X})\,.$}
Thus for any $k\in\mathbb{N}_{0}$ we have
$\frac{1}{MN}\cdot\big|\big\{j\in[MN]\colon\;\|P(\overline{s}_{v_{j}},\,\overline{t})\|=k\big\}\big|=\sum_{\begin{subarray}{c}Y,\,Y\subset X\\ |Y|=k\end{subarray}}\mathrm{Pr}(A_{Y,X})\,.$}
The stated estimate therefore follows from item 1 of Theorem 2.2. $\Box$
3 Concluding remarks
An interesting direction for further research is the following. Suppose that $(P,\overline{S})$, where
$\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,$,
are two $(p,q,n)$-proofreaders and that $\overline{t}\in B_{n}$ is a text. We do not have access to $\overline{t}$ and to $m=\|\overline{t}\|$, but the multiset of triples of values
$V=\{(P(\overline{s}_{v_{j}},\,\overline{t}),\,P(\overline{u}_{w_{j}},\,\overline{t}),\,P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t})))\in\mathbb{N}_{0}^{3}\colon\;j\in[MN]\}$
is known. Can we recover the number of errors $m$, or at least some bounds on $m$, from $V$? The parameter $n$ is of course known but we get two versions of this problem depending on whether $p$ and $q$ are known or not.
Definitions 1.1 and 1.3 with accompanying Propositions 1.2 and 1.4 are probably more important results than Theorem 1.5 whose deduction from the well known Theorem 2.3 is relatively straightforward. The two definitions provide realistic models (RM) of Pólya’s proofreaders. Indeed, if
$\overline{s}=(\overline{s}_{1},\,\overline{s}_{2},\,\dots,\,\overline{s}_{N})$
are runs of a $(p,n)$-proofreader $(P,\overline{s})$ and $\overline{t}\in B_{n}$ is a text, then we can, at least in principle, physically go through all $N=l^{n}$ binary $n$-tuples $\overline{s}_{j}$ and compute all numbers $\|P(\overline{s}_{j},\overline{t})\|$ of errors the proofreader detects in the text during all $N$ runs.
For this to be possible, it is crucial that $p\in\mathbb{Q}$. If $p=1/\sqrt{2}$, say, then no RM for a proofreader with efficiency $p$ exists. For irrational $p$ one has to be content with probabilistic models (PM), of random subsets containing (independently) a given element with probability $p$. Their construction is easy, but they cannot be physically realized in the same sense in which RM can.
Some paradoxes in probability theory arise when a “practical” problem is modeled right away with a PM, while one should start carefully with a RM. The best example of this is the well known Bertrand paradox. This paradox begins as a “practical” problem: what is the probability that a random chord in the unit circle is longer than the length $\sqrt{3}$ of the side of the inscribed equilateral triangle. Solutions start with various PMs, which automatically assume dimensionless points and chords with zero width. This is a mistake, no such objects exist in reality. Instead, one should start with a RM where points have size $s>0$ and chords have width $w>0$. Once such RM is established, one can send $s,w\to 0$ and see what happens in the limit. This is the approach in [5], which we hope to rework and improve in the near future.